If the test statistic is 2.29 for a two-sided t-test for the difference in two means for which n1 = 13, n2 = 17, s1 = 7, and s2 = 5, what's the p-value?
Answer #1

Solution:

We are given

Test statistic = t = 2.29

Test is two tailed.

n1 = 13

n2 = 17

df = n1 + n2 – 2 = 13 + 17 – 2 = 28

So, by using t-table, p-value is given as below:

P-value = 0.0298

(by using t-table or excel)

[Excel command: =TDIST(2.29, 28, 2)]